# Conformal Decision Theory: Safe Autonomous Decisions from Imperfect Predictions

### v1

@@ I Introduction
Autonomous systems increasingly rely on complex learned models to supply predictions that are the basis for decision-making. Self-driving cars rely on deep neural networks [2, 3, 4, 5] to plan paths around nearby pedestrians, robotic manipulators leverage learned grasp models [6] to plan high-throughput pick-and-place maneuvers in factories, and AI-enabled stock trading agents optimize the financial future of investors [7]. There is a conceptual gap, however, between prediction and decision-making, and it remains an open challenge to ensure that systems make good decisions despite imperfect predictions.
One common strategy is to calculate the uncertainty in the predictions independently of their downstream effect on the decision [8, 9, 10, 11, 12]. For example, one can use conformal prediction [13, 14, 15, 16, 17, 18] to form uncertainty sets that cover the ground truth outcomes of all predictions uniformly. Then, the robot can pick any decision that is safe with respect to these sets. This is guaranteed statistically to result in safe autonomous behavior, without any assumption on the underlying distribution or model. This strategy has been used to provide safety assurances in robot navigation and control [19, 20, 21, 22, 23, 24], early warning systems (e.g., collision alerts) [25], out-of-distribution detection [26, 27], probabilistic pose estimation [28], and for large language models [29]. However, this approach decouples prediction uncertainty from decision-making. What if we could solve the problem all-at-once, directly controling decision-making risk, and bypassing the need to construct prediction sets entirely?
@@
The main potential impact of this work is to broaden the scope of conformal prediction. Our methods are more appropriate for disciplines that focus on decision-making, such as control theory, reinforcement learning, and logistics. In these disciplines, algorithms are ultimately evaluated by the decisions, not the predictions, that they make. Furthermore, there are many settings where it does not make sense to construct prediction sets, and our technique can provide a distribution-free outlook for such problems (see, e.g., Section IV-B).
II Conformal Decision Theory
@@
Mathematically, the quality of the decision-making is quantified by a loss function $\mathcal{L}:\mathcal{U}\times\mathcal{Y}\rightarrow[0,1]$. [^1] The framework works for any bounded loss, but we assume the loss to be in $[0,1]$ for simplicity. Often, the loss is more likely to be large when aggressive decisions are taken—i.e., when $\lambda$ is large. For example, $\mathcal{L}$ may be the distance from the planned spline $D^{\lambda}_{t}$ to the nearest human $y_{t}$. Aggressive decisions can be unsafe, but taking $\lambda$ too small yields conservative and under-performing decisions.
We seek an algorithm for adapting $\lambda_{t}$ (and thus the corresponding decision $D^{\lambda}_{t}$) at each time step such that the average loss is controlled in hindsight for any realization of an input-target sequence $\{(x_{t},y_{t})\}^{T}_{t=1}$. This is commonly known as the adversarial sequence model [32, 15]. In this setting, our goal is to set $\lambda_{1:T}$ to achieve a long-term risk bound:
find $\text{find }\lambda_{1:T}\text{ s.t. }\hat{R}_{T}(\mathcal{D},\lambda_{1:T})\leq\varepsilon+\frac{C}{T},$ (2)
where $\varepsilon$ is a pre-defined risk level in $[0,1]$, $C$ is a (small) constant, and
@@
The bound in (2) can be readily extended to
$\hat{R}_{T}(\mathcal{D},\lambda_{1:T})\leq\varepsilon+\frac{C\cdot h(T)}{T},$ (4)
where $h(T)$ is any sublinear function; i.e., one where $h(T)/T\to 0$ as $T\to\infty$.
III Theory & Conformal Controller Algorithm
In this section, we prove the core theoretical results behind Conformal Decision Theory. Specifically, we show that any sequence of families of decision functions $\mathcal{D}_{1:T}$ that are eventually safe can be calibrated online to achieve bounded long-term risk. We then introduce an algorithm called ConformalController which solves Equation (2) under the assumption of eventual safety.
@@ III Theory & Conformal Controller Algorithm
Intuitively, this condition says that there exists a safe value $\lambda^{\rm safe}$ such that if the conformal control variable lands below that value, it will incur a low risk $\varepsilon^{\rm safe}$ after no more than $K$ time steps. For example, even the most conservative robot planner may not be able to change its trajectory fast enough in a single timestep, but it could possibly do so in $K$ time steps. Note that this is a strictly weaker assumption than that used for the proofs in other works, such as [15, 33, 31], which require $K=1$. Conformal controllers are simple yet efficient algorithms that solve the Conformal Decision Theory problem stated in Equation (2). An example is below.
@@
Consider the following update rule for $\lambda$:
$\lambda_{t+1}=\lambda_{t}+\eta\left(\varepsilon-\ell_{t}\right),$ (6)
where $\eta>0$. If $\lambda_{1}\geq\lambda^{\rm safe}-\eta$ and $\mathcal{D}_{1:T}$ satisfies Definition 1 for a given $K\geq 1$ and $\varepsilon^{\rm safe}\leq\varepsilon$, then for any realization of the data, the empirical risk is bounded:
@@
$\lambda_{t}=\lambda_{1}+\eta\sum\limits_{s=1}^{t}(\varepsilon-\ell_{s}).$ (8)
Dividing both sides by $-\eta T$ yields
$\hat{R}_{t}(\lambda_{1:t})=\frac{1}{t}\sum\limits_{s=1}^{t}\ell_{s}$ $=\varepsilon+\frac{\lambda_{1}-\lambda_{t}}{\eta t}.$
To conclude, we just need to show that $\lambda_{t}\geq O(K)$, which is shown in the following Lemma 1.1. ∎
@@
For the sequence in Equation  6, with probability one, we have that the parameter $\lambda_{t}$ is bounded below by $\lambda_{t}\geq\lambda^{\rm safe}-K\eta$, for all $t\in[T]$.
@@
First note that the maximal change in the parameter is $\sup_{t\in[T]}|\lambda_{t+1}-\lambda_{t}|<\eta$, because $\ell_{t}\in[0,1]$ and $\varepsilon\in[0,1]$. We will then proceed by contradiction: Assume that with non-zero probability, $\inf_{u\in[T]}\lambda_{u}<\lambda^{\rm safe}-K\eta$. Denote $t=\arg\min_{u\in[T]}\{u,\lambda_{u}<\lambda^{\rm safe}-K\eta\}$. That is, $t$ is the first instant when the parameter goes below that lower bound. Because the max difference between successive steps is $\eta$, we can prove recursively that $\forall k\in\{0,\dots,K\},\lambda_{t-k}<\lambda^{\rm safe}-(K-k)\eta$. Note that, from those inequalities, we deduce that $t>K$ since $\lambda_{1}\geq\lambda^{\rm safe}-\eta$. By recursively applying the update rule $\lambda_{t}=\lambda_{t-K}+K\eta(\varepsilon-\frac{1}{K}\sum_{k=1}^{K}\ell_{t-k})$, we have:
@@
$\Longrightarrow\frac{1}{K}\sum_{k=1}^{K}\ell_{t-k}\leq\varepsilon^{\rm safe}$
$\Longrightarrow\lambda_{t}=\lambda_{t-K}+K\eta\left(\varepsilon-\frac{1}{K}\sum_{k=1}^{K}\ell_{t-k}\right)\geq\lambda_{t-K}+K\eta(\varepsilon-\varepsilon^{\rm safe})$
@@
Conformal decision theory can also be applied in the so-called batch setting, wherein a separate calibration dataset is available for learning a safe decision. Here, a dataset or simulator allows for offline experimentation to quantify the risk of different decisions, e.g., offline RL. This requires a different statistical setup.
Consider the case of $n+1$ exchangeable decision functions $D_{1}(\lambda),\ldots,D_{n+1}(\lambda)$ and an associated loss function $\mathcal{L}$ taking a decision and returning a value in $[0,1]$. The first $n$ decision functions will be used for calibration of a parameter $\hat{\lambda}$ that will be used in the final decision. These exchangeable decision functions may be produced, for example, by applying a single decision function to a sequence of exchangeable data points. For the sake of simplicity, we assume that the decisions have monotone loss, i.e., that for all $i$,
@@
Though the validity of the algorithm follows from the theory of conformal risk control, it is substantially different in practice and deserves further study. Specifically, unlike the previous methods, in order to calculate $\hat{\lambda}$, one must iterate through a sequence of counterfactual decisions (in other words, putative values of $\lambda$) and test what the effect would have been. This restricts the applications of the batch algorithm and also presents an opportunity for future work to make it more efficient and expand its scope.
IV Experiments
@@
IV-A Robot Navigation in Stanford Drone Dataset
Robot navigation around people must balance safety (i.e., not colliding with humans) and efficiency (i.e., the robot makes progress towards a goal). To ensure that the risk of collision is low while still making progress to the goal, the robot will calibrate its cost function at run-time using a conformal controller.
@@ IV-A Robot Navigation in Stanford Drone Dataset
where $u^{\mathrm{pos}}_{t}$ is the robot’s current position. To make this value bounded, we clip the loss to the size of the video.
Metrics. We measure a boolean safe variable indicating if the robot did not ever collide with a human. We also measure a boolean success variable if the robot reached the goal location by the end of the interaction episode (i.e., length of video in the SDD). We also measure the time to reach the goal location and the minimum, mean, and $\{5\%,10\%,25\%,50\%\}$ quantiles of the distance to the nearest human.
@@
Figure 3: Stanford Drone Dataset. (Top) Trajectories of $\lambda_{t}$ (calibrated by CC) and $\alpha_{t}$ (used by ACI to calibrate prediction sets). When $\alpha_{t}\leq 0$, ACI returns infinite set and the robot stops. (Bottom) Distance to the nearest human over time. $\lambda_{t}$ is large when the robot is close to a human, while $\alpha_{t}$ is unrelated. The $\lambda_{t}$ trajectory is shorter because it reaches the goal faster.
IV-B Manufacturing Assembly Line Robot
@@
IV-C Stock Trading Agent
@@
V Discussion & Conclusion
In this paper, we introduce Conformal Decision Theory, a theoretical and algorithmic framework for producing safe decisions despite being based on imperfect machine-learning predictions. We have described our method in both the online, adversarial setting, and also the batch, exchangeable setting. The main difference between the two is that the online algorithms we present are computationally trivial, while the batch setting can require evaluating a large amount of counterfactual decisions (indexed by different choices of $\lambda$) on every calibration point. Though this can be done with binary search, it still presents operational challenges. One path for future work may be to test the method in settings where simulators or data sets can support this form of offline policy evaluation. Another may be to develop formally valid approximations of the batch technique which preserve risk control while being more practical. Furthermore, extensions of the batch technique to non-exchangeable settings are readily available, e.g., by use of the techniques in [38], and could be evaluated. Finally, future work may additionally consider optimizing the conformal control variable to maximize utility, perhaps also subject to the constraint of risk control, bringing the work closer to the classical statistical decision theory of Lehmann [39], von Neuman and Morgenstern [40], and others.

### v3

@@ I Introduction
Autonomous systems increasingly rely on complex learned models to supply predictions that are the basis for decision-making. Self-driving cars rely on deep neural networks [2, 3, 4, 5] to plan paths around nearby pedestrians, robotic manipulators leverage learned grasp models [6] to plan high-throughput pick-and-place maneuvers in factories, and AI-enabled trading agents optimize the financial future of investors [7]. There is a conceptual gap between prediction and decision-making, and it remains a challenge to ensure that systems make good decisions despite imperfect predictions.
One increasingly popular strategy is to quantify the uncertainty in the predictions independently of their downstream effect on the decision via conformal prediction (CP) [8, 9, 10, 11, 12, 13]. This approach has become popular because, when used to provide simultaneous prediction sets on all outcomes, conformal prediction provides statistical guarantees of safe autonomous behavior without any assumption on the underlying distribution or model. This application of CP has shown impact in robot navigation [14, 15, 16, 17], early warning systems [18], out-of-distribution detection [19, 20], probabilistic pose estimation [21], and for large language models [22]. However, the requirement of simultaneous coverage is challenging to satisfy and for many decision systems is excessive. What if we could provide statistical guarantees, as in CP, directly on our decisions, bypassing the need to construct prediction sets?
@@
The main potential impact of this work is to broaden the scope of conformal prediction. Our methods are more appropriate for disciplines that focus on decision-making, such as control theory, reinforcement learning, and logistics. In these disciplines, algorithms are ultimately evaluated by the decisions, not the predictions, that they make. Furthermore, there are many settings where it does not make sense to construct prediction sets, and our technique can provide a distribution-free outlook for such problems (see, e.g., Section V-B).
II Related Work
Decision-Making Under Predictive Uncertainty. Within the machine learning and statistics community, uncertainty quantification of prediction models has been studied widely, from conformal prediction to Bayesian neural networks to ensembles [25, 26, 27, 28, 29]. Instead of focusing on prediction calibration alone, the controls and optimization community have coupled prediction uncertainty with safe (i.e., risk bounded) decision-making via chance-constrained optimal control [30, 31] and scenario optimization [32, 33]. The former typically constructs prediction sets that are used as constraints while the latter safeguards against samples drawn from the prediction model. Instead of directly calibrating the output of upstream prediction modules or solving decision-making problems under probabilistic constraints, this work presents a theoretical and algorithmic approach to tuning the robot’s decision risk directly as a function of historical decision-making performance.
Online Learning & Nonstochastic Control. The method herein is reminiscent of online learning, and specifically the online gradient descent (OGD) update of [34]. The connection is most apparent when examining the forthcoming Equation (5) with $\ell_{t}=\mathbf{1}\{y_{t}\in\mathcal{C}(x_{t})\}$; this recovers the ACI algorithm of [10], which is OGD on the quantile loss [29]. However, the update in (5) is substantially more general because it incorporates arbitrary decision rules, and reframing it as OGD on an analytic loss function is generally impossible. Furthermore, the guarantees in [10] are not to our knowledge recoverable by existing regret analyses from online convex optimization and nonstochastic control, e.g.,  [35, 36, 37]. However, the guarantees do share a retrospective flavor, in that, like regret analyses, they provide guarantees on average over the observed history.
III Conformal Decision Theory
@@
Mathematically, the quality of the decision-making is quantified by a loss function $\mathcal{L}:\mathcal{U}\times\mathcal{Y}\rightarrow[0,1]$. [^1] The framework works for any bounded loss, but we assume the loss to be in $[0,1]$ for simplicity. Often, the loss is more likely to be large when aggressive decisions are taken—i.e., when $\lambda$ is large. Aggressive decisions can be unsafe, but taking $\lambda$ too small yields conservative and under-performing decisions.
We seek an algorithm for adapting $\lambda_{t}$ (and thus the corresponding decision $D^{\lambda}_{t}$) at each time step such that the average loss is controlled in hindsight for any realization of an input-target sequence $\{(x_{t},y_{t})\}^{T}_{t=1}$. This is commonly known as the adversarial sequence model [38, 10]. Here, our goal is to set $\lambda_{1:T}$ to achieve a long-term risk bound:
find $\text{find }\lambda_{1:T}\text{ s.t. }\hat{R}_{T}(\mathcal{D}_{1:T},\lambda_{1:T})\leq\varepsilon+\frac{C\cdot h(T)}{T},$ (2)
where $\varepsilon$ is a pre-defined risk level in $[0,1]$, $C$ is a (small) constant, $h(T)$ is any sublinear function; i.e., one where $h(T)/T\to 0$ as $T\to\infty$, and
@@
We will omit $\mathcal{D}_{1:T}$ in the notation of risk when the sequence of family of decision functions is clear from the context
IV Theory & Conformal Controller Algorithm
In this section, we prove the core theoretical results behind Conformal Decision Theory. Specifically, we show that any sequence of families of decision functions $\mathcal{D}_{1:T}$ that are eventually safe can be calibrated online to achieve bounded long-term risk. We then introduce an example of conformal controller which solves Equation (2) under the assumption of eventual safety.
@@ IV Theory & Conformal Controller Algorithm
Intuitively, this condition says that there exists a safe value $\lambda^{\rm safe}$ such that if the conformal control variable lands below that value, it will incur a low risk $\varepsilon^{\rm safe}$ after no more than $K$ time steps. For example, even the most conservative robot planner may not be able to change its trajectory fast enough in a single timestep, but it could possibly do so in $K$ time steps. For general decision-making, the existence of a safe decision function is not guaranteed, and requires domain-specific knowledge (e.g., when the loss function captures the distance between agents [39, 40, 41]). But when the decision is a prediction set, a safe decision function is trivial because you can always output the entire space. Note that the eventually safe is a strictly weaker assumption than that used for the proofs in other works, such as [10, 42, 24], which require $K=1$. Moreover, conformal controllers are simple yet efficient algorithms that solve the Conformal Decision Theory problem stated in Equation (2). An example is below.
@@
Consider the following update rule for $\lambda_{1:T}$:
$\lambda_{t+1}=\lambda_{t}+\eta\left(\varepsilon-\ell_{t}\right),\forall t\in[T]$ (5)
where $\eta>0$ and $\ell_{t}:=\mathcal{L}(D_{t}^{\lambda_{t}}(x_{t}),y_{t})$.
If $\lambda_{1}\geq\lambda^{\rm safe}-\eta$ and $\mathcal{D}_{1:T}$ satisfies Definition 1 for a given $K\geq 1$ and $\varepsilon^{\rm safe}\leq\varepsilon$, then for any realization of the data, the empirical risk is bounded:
@@
$\lambda_{t+1}=\lambda_{1}+\eta\sum\limits_{s=1}^{t}(\varepsilon-\ell_{s}).$ (7)
By isolating $\sum_{s=1}^{t}\ell_{s}$ on one side and moving all other terms to the right-hand side, we obtain:
$\hat{R}_{t}(\lambda_{1:t})=\frac{1}{t}\sum\limits_{s=1}^{t}\ell_{s}$ $=\varepsilon+\frac{\lambda_{1}-\lambda_{t+1}}{\eta t}.$
To conclude, we just need to lower bound $\lambda_{t}$ by a constant w.r.t $t$ which is done in the following Lemma 1.1. ∎
@@
For the sequence in Equation  5, with $\lambda_{1}\geq\lambda^{\rm safe}-\eta$ we have that the parameter $\lambda_{t}$ is bounded below by $\lambda_{t}\geq\lambda^{\rm safe}-K\eta$, for all $t\in[T+1]$.
@@
First note that the maximal change in the parameter is $\sup_{s\in[T]}|\lambda_{s+1}-\lambda_{s}|<\eta$, because $\ell_{s}\in[0,1]$ and $\varepsilon\in[0,1]$. We will then proceed by contradiction: Assume that $\inf_{s\in[T+1]}\lambda_{s}<\lambda^{\rm safe}-K\eta$. Denote $t=\arg\min_{s\in[T+1]}\{s:\lambda_{s}<\lambda^{\rm safe}-K\eta\}$. That is, $t$ is the first instant when the parameter goes below that lower bound. Then, by definition of $t$, $\forall s<t,\lambda_{t}<\lambda^{\rm safe}-K\eta\leq\lambda_{s}$.
Because the max difference between successive steps is $\eta$, we can prove recursively that $\forall k\in\{0,\dots,K\},\lambda_{t-k}<\lambda^{\rm safe}-(K-k)\eta$. Note that, from those inequalities, we deduce that $t>K$ since $\lambda_{1}\geq\lambda^{\rm safe}-\eta$. By recursively applying the update rule $\lambda_{t}=\lambda_{t-K}+K\eta(\varepsilon-\frac{1}{K}\sum_{k=1}^{K}\ell_{t-k})$, we have:
@@
[…]
$\Longrightarrow\lambda_{t}=\lambda_{t-K}+K\eta\left(\varepsilon-\frac{1}{K}\sum_{k=1}^{K}\ell_{t-k}\right)$
$\Longrightarrow\lambda_{t}\geq\lambda_{t-K}+K\eta(\varepsilon-\varepsilon^{\rm safe})$
@@
Remark 1.
The assumption $\lambda_{1}\geq\lambda^{\rm safe}-\eta$ is not necessary to prove that $\hat{R}_{t}(\lambda_{1:t})\leq\varepsilon+O(1/t)$. Intuitively, two scenarios can occur:
  1. 1.
    If $\forall k\in[K],\lambda_{k}\leq\lambda^{\rm safe}$, then the empirical risk over the first $K$ steps will be upper bounded by $\varepsilon^{\rm safe}$. In this case, we need only to upper bound the risk between $K+1$ and $T$, which can be achieved using the previous theorem or this remark.
  2. 2.
    If there exists a $k\in[K]$ such that $\lambda_{k}>\lambda^{\rm safe}$, we can apply the previous theorem to upper bound the risk between $k$ and $T+1$. The cumulative loss between $1$ and $k$ is upper bounded by $k$, which is $o(1)$.
@@
Conformal decision theory can also be applied in the so-called batch setting, wherein a separate calibration dataset is available for learning a safe decision. Here, a dataset or simulator allows for offline experimentation to quantify the risk of different decisions, e.g., offline RL. This requires a different statistical setup. Consider the case of $n+1$ exchangeable decision functions $D_{1}(\lambda),\ldots,D_{n+1}(\lambda)$ and an associated loss function $\mathcal{L}$ taking a decision and returning a value in $[0,1]$. The first $n$ decision functions will be used for calibration of a parameter $\hat{\lambda}$ that will be used in the final decision. These exchangeable decision functions may be produced, for example, by applying a single decision function to a sequence of exchangeable data points. For the sake of simplicity, we assume that the decisions have monotone loss, i.e., that for all $i$,
@@
Though the validity of the algorithm follows from the theory of conformal risk control, it is substantially different in practice and deserves further study. Specifically, unlike the previous methods, in order to calculate $\hat{\lambda}$, one must iterate through a sequence of counterfactual decisions (possible values of $\lambda$) and evaluate what the loss would have been. This restricts the applications of the batch algorithm and also presents an opportunity for future work to make it more efficient and expand its scope.
V Experiments
@@
V-A Robot Navigation in Stanford Drone Dataset
Robot navigation around people must balance safety (i.e., not colliding with humans) and efficiency (i.e., the robot makes progress towards a goal). To ensure that the risk of collision is low while still making progress to the goal, the robot will calibrate its cost function at run-time using a conformal controller (CC).
@@ V-A Robot Navigation in Stanford Drone Dataset
where $u^{\mathrm{pos}}_{t}$ is the robot’s current position. To make this value bounded, we clip the loss to the size of the video. Note that because we use a negative loss, we also changed $\lambda$ so that the larger $\lambda$, the more conservative the decision.
Metrics. We measure a boolean safe variable indicating if the robot did not ever collide with a human. We also measure a boolean success variable if the robot reached the goal location by the end of the interaction episode (i.e., length of video in the dataset). We also measure the time to reach the goal location and the minimum, mean, and $\{5\%,10\%,25\%,50\%\}$ quantiles of the distance to the nearest human.
Figure 2: Stanford Drone Dataset. (Top) Trajectories of $\lambda_{t}$ (calibrated by CC) and $\alpha_{t}$ (from ACI to calibrate sets). When $\alpha_{t}\leq 0$, ACI returns infinite set and the robot stops. (Bottom) Distance to the nearest human over time. $\lambda_{t}$ is large when the robot is close to human, while $\alpha_{t}$ is unrelated. The $\lambda_{t}$ trajectory is shorter because it reaches the goal faster.
V-B Manufacturing Assembly Line Robot
@@
V-C Stock Trading Agent
@@
VI Discussion & Conclusion
In this paper, we introduce Conformal Decision Theory, a theoretical and algorithmic framework for producing safe decisions despite being based on imperfect machine-learning predictions. We have described our method in both the online adversarial setting, and also the batch exchangeable setting. The main difference between the two is that the online algorithms we present are computationally trivial, while the batch setting can require evaluating a large amount of counterfactual decisions (indexed by different choices of $\lambda$) on every calibration point. Though this can be done with binary search, it still presents operational challenges. One path for future work may be to test the method in settings where simulators or data sets can support this form of offline policy evaluation. Another may be to develop formally valid approximations of the batch technique which preserve risk control while being more practical. Furthermore, extensions of the batch technique to non-exchangeable settings are readily available, e.g., by use of the techniques in [47], and could be evaluated.
Finally, despite $\lambda$ being 1-dimensional, our procedure can index an arbitrary set of decisions. Consider a set of decisions $\mathcal{D}$, a utility predictor $\hat{u}(d;x)$ where $d\in\mathcal{D}$, and a loss predictor ^$\hat{\mathcal{L}}(d;x)$, we can maximize utility subject to the constraint that our predicted loss is controlled:
$D_{t}=\arg\max_{d\in\mathcal{D}}\quad$ $\hat{u}(d;x_{t})$
s.t. ^$\hat{\mathcal{L}}(d;x_{t})\leq\lambda_{t}.$
This will work as long as we revert to a safe decision if $\lambda_{t}\leq\lambda^{\rm safe}$; where the sequence $\lambda_{t}$ is defined in Equation 5. However, no guarantees on utility are provided. This topic would be a great avenue for future work, bringing conformal prediction closer to the classical statistical decision theory of Lehmann [48], von Neuman and Morgenstern [49], and others.